how do I put this quadratic equation into general form? The fraction is making it complicated. Thank you.
1/5(3x^2-10)=12x

Answer by ankor@dixie-net.com(22740)

1/5(3x^2-10) = 12x
multiply both sides by 5
3x^2 - 10 = 60x
3x^2 - 60x - 10 = 0